OpenStudy (anonymous):

how much heat is released when 0.5kg of steam at 100◦C turns to order at a temperature at 20◦C ?

OpenStudy (ghazi):

Q=m*c*(T2-T1) use c= specific heat capacity of water= 4.18, m= 0.5kg and T2-T1=80

OpenStudy (anonymous):

m = m c t m= 0.5 kg c= 4200 j/kg/k t= 80 put values get answer